OpenStudy (anonymous):

(Is there a search function here?) I'm having trouble wrapping my head around 1A-4, but I think I found the problem. First of all, the actual exercise version has a typo: the last term should have a minus between the numerator terms. The solution writes this correct, but gets G(-x) wrong as far as I can see. It should read: \[G(-x)=\frac{f(-x)-f(x)}{2}=-\frac{f(x)-f(-x)}{2}=-G(x)\]

OpenStudy (phi):

Yes, you are correct, and they muddled both the question and the answer. https://en.wikipedia.org/wiki/Even_and_odd_functions has some simple examples. If they start with the correct expression: \[ f(x) = \frac{ f(x) + f(-x)}{2} + \frac{ f(x) - f(-x)}{2} \] It is easy to add the fractions on the right-hand side to get \[ f(x) = \frac{ f(x) + f(-x)+f(x) - f(-x)}{2} = \frac{2f(x)}{2}= f(x)\] showing the identity is correct. If we define the first fraction as \[ F(x) = \frac{ f(x) + f(-x)}{2} \] we see \[ F(-x) = \frac{ f(-x) + f(-(-x))}{2} = \frac{ f(-x) + f(x)}{2} = F(x)\] which shows F(x) is even similarly, define the second fraction as \[ G(x) = \frac{ f(x) - f(-x)}{2} \] and we find \[ G(-x) = \frac{ f(-x) - f(x)}{2} = - \frac{ f(x) - f(-x)}{2} = -G(x)\] which shows G(x) is odd. (See wikipedia link for details on even/odd )

OpenStudy (anonymous):

I struggled with this for a while until I realized that they had messed up the question.